Use a direct proof to show that the sum of two odd integers(k and l) is even.


1
Expert's answer
2022-06-13T12:03:23-0400

Proof.

Let "k" and "l" be the odd integers. Then there exist some integers "k_0" and "l_0" such that

"k=2k_0+1,"

"l=2l_0+1."

We can find the sum:

"k+l=(2k_0+1)+(2l_0+1)=2k_0+2l_0+2=2(k_0+l_0+1),"

where "k_0+l_0+1" is an integer as a sum of integers.

So "k+l" is divisible by 2 and is even.

The proof is finished.